# The Ginsburg–Sands theorem and computability theory

### v1

@@ 7.2.2. Forcing notion for $L_{G}$
in which each condition is a constructible extension of the previous and for every $\sigma\in T_{F^{i}_{p_{n},b^{i}_{n}},V^{i},\emptyset}$ of maximal length there exist $x_{1}>x_{0}>k$ such that $\Phi^{\sigma}_{e}(x_{0})\operatorname{\downarrow}=\Phi^{\sigma}_{e}(x_{1})\operatorname{\downarrow}=1$. Note that since $P\gg L^{\prime}$ and $L\gg X\oplus Y$, it is possible for $P$ to search through all indices of infinite $\Pi^{0}_{1}(X\oplus Y)$ classes. (For this, it suffices just that $P\geq_{\text{T}}(X\oplus Y)^{\prime}$.) Hence, given $k$, it follows by Lemmas 7.7 and 7.8 that $P$ can uniformly search for a sequence of preconditions as above. By assumption, this search must always succeed. Let $(p_{n},b^{0}_{n},b^{1}_{n})$ be the last condition in this sequence. Then $P$ can find all the $x_{1}>x_{0}>k$ as above, across all $\sigma\in T_{F^{i}_{p_{n},b^{i}_{n}},V^{i},\emptyset}$, and in particular can find the minimum $x^{k}_{0}$ of all such $x_{0}$, and the maximum $x^{k}_{1}$ of all such $x_{1}$. Thus, $x^{k}_{1}>x^{k}_{0}>k$. Since $\overline{S}$ is $P$-hyperimmune, there must be a $k$ such that $[x^{k}_{0},x^{k}_{1})\subseteq S$. Fix this $k$, and fix the corresponding sequence $(p,b^{0},b^{1})=(p_{0},b^{0}_{0},b^{1}_{0}),\ldots,(p_{n},b^{0}_{n},b^{1}_{n})$ found by $P$ in the above search. Then $(p_{n},b^{0}_{n},b^{1}_{n},V^{0},V^{1})$ forces $\Phi^{P_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $\overline{S}$ on the $i$ side, so we can take our sequence witnessing (2) to be $(p_{0},b^{0}_{0},b^{1}_{0},V^{0},V^{1}),\ldots,(p_{n},b^{0}_{n},b^{1}_{n},V^{0},V^{1})$. ∎
@@ 7.2.4. Putting it all together
Suppose the result is false for $i=0$. First, suppose that there is an $e_{0}$ such that no $p_{m}$ forces $e_{0}\in L_{\mathtt{G}}^{\prime}$ or $e_{0}\notin L_{\mathtt{G}}^{\prime}$ on the 00 side. Fix $e\in\omega$. Then the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+2$ (which appeals to Lemma 7.20 with $e_{0},e$) must force $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $1$ side, while the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+4$ (which appeals to Lemma 7.22 with $e_{0},e$) must force $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the $1$ side. Suppose instead that there is an $e_{0}$ such that no $p_{m}$ forces $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the 00 side. Then the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+3$ (which appeals to Lemma 7.21 with $e_{0},e$) must force $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the $1$ side, while the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+5$ (which appeals to Lemma 7.23 with $e_{0},e$) must force $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $1$ side. ∎
@@ 8. Summary and questions
[…]
Figure 2. The hierarchy of subset principles, with $n\geq 3$ arbitrary. Arrows indicate implications over $\mathsf{RCA}_{0}$; double arrows are strict.
@@
[…]
@@
[…]
@@
[…]
@@
The hierarchy of $\Gamma\text{-}\mathsf{Subset}$ principles is summarized in Figure 2.

### v2

@@ 7.2.2. Forcing notion for $L_{G}$
in which each condition is a constructible extension of the previous and for every $\sigma\in T_{F^{i}_{p_{n},b^{i}_{n}},V^{i},\emptyset}$ of maximal length there exist $x_{1}>x_{0}>k$ such that $\Phi^{\sigma}_{e}(x_{0})\operatorname{\downarrow}=\Phi^{\sigma}_{e}(x_{1})\operatorname{\downarrow}=1$.
Note that since $P\gg L^{\prime}$ and $L\gg X\oplus Y$, it is possible for $P$ to search through all indices of infinite $\Pi^{0}_{1}(X\oplus Y)$ classes. (For this, it suffices just that $P\geq_{\text{T}}(X\oplus Y)^{\prime}$.) Hence, given $k$, it follows by Lemmas 7.7 and 7.8 that $P$ can uniformly search for a sequence of preconditions as above. By assumption, this search must always succeed. Let $(p_{n},b^{0}_{n},b^{1}_{n})$ be the last condition in this sequence. Then $P$ can find all the $x_{1}>x_{0}>k$ as above, across all $\sigma\in T_{F^{i}_{p_{n},b^{i}_{n}},V^{i},\emptyset}$, and in particular can find the minimum $x^{k}_{0}$ of all such $x_{0}$, and the maximum $x^{k}_{1}$ of all such $x_{1}$. Thus, $x^{k}_{1}>x^{k}_{0}>k$.
Since $\overline{S}$ is $P$-hyperimmune, there must be a $k$ such that $[x^{k}_{0},x^{k}_{1})\subseteq S$. Fix this $k$, and fix the corresponding sequence $(p,b^{0},b^{1})=(p_{0},b^{0}_{0},b^{1}_{0}),\ldots,(p_{n},b^{0}_{n},b^{1}_{n})$ found by $P$ in the above search. Then $(p_{n},b^{0}_{n},b^{1}_{n},V^{0},V^{1})$ forces $\Phi^{P_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $\overline{S}$ on the $i$ side, so we can take our sequence witnessing (2) to be $(p_{0},b^{0}_{0},b^{1}_{0},V^{0},V^{1}),\ldots,(p_{n},b^{0}_{n},b^{1}_{n},V^{0},V^{1})$. ∎
@@ 7.2.4. Putting it all together
Suppose the result is false for $i=0$. First, suppose that there is an $e_{0}$ such that no $p_{m}$ forces $e_{0}\in L_{\mathtt{G}}^{\prime}$ or $e_{0}\notin L_{\mathtt{G}}^{\prime}$ on the 00 side. Fix $e\in\omega$. Then the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+2$ (which appeals to Lemma 7.20 with $e_{0},e$) must force $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $1$ side, while the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+4$ (which appeals to Lemma 7.22 with $e_{0},e$) must force $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the $1$ side.
Suppose instead that there is an $e_{0}$ such that no $p_{m}$ forces $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the 00 side. Then the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+3$ (which appeals to Lemma 7.21 with $e_{0},e$) must force $\Phi^{L_{\mathtt{G}}}_{e}$ is not a non-hyperimmunity witness for $S$, or $\Phi^{L_{\mathtt{G}}}_{e}$ is finite, on the $1$ side, while the condition $p_{n_{s}}$ defined at stage $s=8\langle e_{0},e\rangle+5$ (which appeals to Lemma 7.23 with $e_{0},e$) must force $e\in L_{\mathtt{G}}^{\prime}$ or $e\notin L_{\mathtt{G}}^{\prime}$ on the $1$ side. ∎
@@ 8. Summary and questions
${\Delta^{0}_{n+1}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Sigma^{0}_{n}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Delta^{0}_{n}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Sigma^{0}_{3}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Delta^{0}_{3}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathsf{SGST}_{1}}$${\Sigma^{0}_{2}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathsf{D}^{2}_{2}}$${\Delta^{0}_{2}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Delta^{0}_{1}\text{-}\mathsf{Subset}}$${\Sigma^{0}_{1}\text{-}\mathsf{Subset}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathsf{RCA}_{0}}$
Figure 2. The hierarchy of subset principles, with $n\geq 4$ arbitrary. Arrows indicate implications over $\mathsf{RCA}_{0}$; double arrows are strict; interrupted arrows are non-implications.
@@
With regard to Corollary 6.15, which employed a modification of the second jump control proof from [2] to show that $\mathsf{GST}_{1}$ admits low2 solutions, we can ask whether the same result can also be obtained by first jump control. More precisely, in [2, Section 4] this method was employed to show that, given any $P\gg\emptyset^{\prime}$ and any computable $c:[\omega]^{2}\to 2$, there exists an infinite homogeneous set $H$ for $c$ satisfying $H^{\prime}\leq_{\text{T}}P$.
@@
The second jump control proof in [2] was also used to establish a number of conservativity results, perhaps most notably that $\mathsf{RCA}_{0}+\mathsf{I}\Sigma^{0}_{2}+\mathsf{RT}^{2}_{2}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{I}\Sigma^{0}_{2}$ ([2], Theorem 10.2). This had broadly the same outline as the low2 proof, but differed in one important respect. Recall that this proof only required $A$ to be $\Delta^{0}_{2}$ (actually, $\Sigma^{0}_{2}$) when building a subset of $A$; it worked equally well for $\Delta^{0}_{3}$ (and hence in particular, $\Pi^{0}_{2}$) sets when building a subset of $\overline{A}$. This is what made it possible for us to adapt the proof and obtain Theorem 6.14. The conservativity proof, by contrast, requires $A$ to be $\Delta^{0}_{2}$ for both parts, and thus does not similarly lift to $\Sigma^{0}_{2}$ sets.
@@
Finally, we can ask some more general questions about the $\Gamma\text{-}\mathsf{Subset}$ principles. By Theorem 7.1, $\Delta^{0}_{2}\text{-}\mathsf{Subset}$ does not imply $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$, but our proof does not relativize to show that $\Delta^{0}_{n}\text{-}\mathsf{Subset}$ does not imply $\Sigma^{0}_{n}\text{-}\mathsf{Subset}$ for any higher $n$.
@@
Is there an $n\geq 3$ such that $\Delta^{0}_{n}\text{-}\mathsf{Subset}$ implies $\Sigma^{0}_{n}\text{-}\mathsf{Subset}$?
It is also true that $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$ does not imply $\Delta^{0}_{3}\text{-}\mathsf{Subset}$. Indeed, by relativizing the main result of Downey, Hirschfeldt, Lempp, and Solomon [12] it follows that for all $n\geq 2$ there is a $\Delta^{0}_{n}$ set with no infinite subset in it or its complement that is low over $\emptyset^{(n-2)}$ (see also the discussion in [12, p. 1372]). Taking $n=3$ implies, in particular, that there is a $\Delta^{0}_{3}$ set with no infinite low2 subset in it or its complement. Hence, $\Delta^{0}_{3}\text{-}\mathsf{Subset}$ fails in the model of $\mathsf{GST}_{1}$ (and also $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$) constructed in Corollary 6.16 above. Again, we do not know if this holds more generally.
Question 8.8.
@@
Monin and Patey [37, Theorem 1.4] have shown that for every $n\geq 2$, every $\Delta^{0}_{n}$ set has an infinite lown subset in it or its complement. Combined with the aforementioned result from [12], this implies that $\Delta^{0}_{n}\text{-}\mathsf{Subset}$ does not imply $\Delta^{0}_{n+1}\text{-}\mathsf{Subset}$ for all $n\geq 2$. Thus, we at least know that 8.7 and 8.8 cannot both be answered affirmatively for the same $n$. The hierarchy of $\Gamma\text{-}\mathsf{Subset}$ principles is summarized in Figure 2.